anyone have more teas exam question ?
Answer #1

3. 1. Which is an anterior muscle?
a. gluteus maximas
b. biceps femoris
c. adductor magnus
d. quadriceps femoris

"d"
Anterior muscle is located near the front of the body. Of those listed, the quadriceps is the most anterior

2. The lateral side of the right knee would be:
a. the kneecap
b. closest to the left knee
c. farthest from the left knee
d. on the underside of the knee

"c" Something that is lateral is toward the outer side of the body. The side of the right knee farthest from th left knee would fit this description

3. The ovaries are part of the ________system.

a. skeletal
b. nervous
c. lymphatic
d. reproductive

"d" The ovaries are the egg-producing reproductive organs

4. In a whiplash injury, you might expect to see:

a. cardiopulmonary problems
b. side-to-side spinal curvature
c. eventual injury to ligaments
d. traumatic injury to ligaments

"d" Because of the hyperflexion and hyperextension caused by such a collision, it is not uncommon to find tears or stretching of ligaments of the spine. The injury is front to back

5. Which is an example of deductive reasoning?

A. Ten of the 100 applicants are left-handed; therefore 10% of the general population must be left-handed
B. I know several left-handed artists; therefore left-handed people must have artistic aptitude
C. Left-handed people are artistic; Jon is left-handed. Therefore, John must be artistic.
D. I see that Jon is left-handed. Jon is a gifted artist. Therefore left-handed people must be gifted artists.

"C" In deductive reasoning, a hypothesis follows from a set of premises. The premises may be true, in which case the reasoning is sound, or they may be untrue or at least dubious, as in this example.

6. If gas A has four times the molar mass of gas B, you would expect it to diffuse through a plug:

a. at half the rate of gas B
b. at twice the rate of gas B
c. at a quarter the rate of gas B
d. at four times the rate of gas B

"a" Set up as a simple equation if it helps - Suppose that gas A has a molar mass of M1 and gas B has a molar mass four times that M2. Now suppose that that M1 = 1, and M2 = 4. THe square root of 4/1 = 2/1 so the rate of effusion for gas A is half that of gas B

7. What is the percent by volume of ethanol in 120ml of a solution containing 35ml of ethanol dissolved in water?

a. 22.6%
b. 29.2 %
c. 35%
d. 41.1%

"b" Essentially you are determining the concentration as if it were in a 100ml solution, which is what percent really means. You could set up the equation:

35/120 = x/100
3500/100x
x = 29.2

8. Three students measured the mass of a product of combustion. They recorded measurements of 5.14g, 5.16g, and 5.17g. If the known mass of the product is 5.3g, the students measurements were _____________?

a. accurate
b. precise
c. both precise and accurate
d. neither accurate or precise

"b" The students measurements were precise in that they correspond closely with one another. However, since the true measurement differed from theirs by at least .13g, the measurements were not terribly accurate.

9. Which type of nutrient does not provide the body with energy?

a. vitamin
b. carbohydrate
c. fat
d. protein

"a" Some nutrients supply energy, others support metabolism. Nutrients in the first set include carbohydrates, proteins, and fats. Those in the second set include water, vitamins, minerals.

10. Why might certain young people be underweight?

a. They are growing in height faster than they are gaining weigh
b. they eat little protein and too much fat and sugar
c. they eat several meals a day and fail to exercise enough
d. they seldom eat a meal that contains foods from all food groups

"a" This is a common cause of being underweight in adolescents